find the area of a circle with a radius of 4 1/2 inches. Round your answer to the nearest hundredth
Answer by nyc_function(2741)

A = pi(r^2)

4 1/2 can be written as an improper fraction 9/2.

So, let r = 9/2

A = pi(9/2)^2

Let pi = 3.14

A = 3.14(9/2)^2

A = 3.14(81/4)

A = 3.14(20.25)

A = 63.585in^2

We now round to the nearest hundredths AKA second decimal place.

A = 63.59in^2